The raw database from which the catalogs are derived can be examined in /mit/mitsfs/Public/datadex on Athena. There is also a Web interface at http://www.mit.edu:8001/pinkdex which can be used for title and author searches. Our homepage is at http://www.mit.edu/∼mitsfs/. Please mention errors in the catalog on the other side of this page. New first copies of books may be on the New Books shelf rather than with the rest of their category. Monthly catalog supplements of new first copies are kept at the New Books shelf for a year. The CODE column entries look like (code):(multiplicity), or just (code) if we’ve only one. The codes are defined as follows (ask a Keyholder what the categories are):

C/P, C/PA  Circulating Paperback, Circulating Paperback Anthology
P, PA  [Reserve] Paperback, Paperback Anthology
H, HA  [Circulating] Hardcover, Hardcover Anthology
LP, LPA  [Reserve] Large P, PA
VLP, VLPA  [Reserve] Very Large P, PA
R/REF-H, R/REF-P  Reserve Reference H, P (non-fiction)
C/REF-H, C/REF-P  Circulating Reference H, P (non-fiction)
R/IND, R/IND-F  Reserve Index, Index Folder
SRL/IND  Special Reserve Locked Index
BS  [Reserve] Binary Star Double
D  [Reserve] Ace Double
GRD, GZD, OVD, TD  [Reserve] Gryphon, Gollancz, Ocean View, Tor Double
C/BS, C/D, C/TD, ...  Circulating Doubles
GN  [Reserve] Galaxy Novel
SFWA  [Reserve] SFWA Hardcover
SFWA-P, SFWA-VLP  [Reserve] SFWA P, VLP
SFWA-TD  [Reserve] SFWA Tor Double
SR-H, SR-P, SR-PA, ...  Special Reserve H, P, PA, ...
SR-VLH, SR-VLHA  Special Reserve Very Large H, HA
SR-DNESFA  Special Reserve NESFA Double
SR-MM  Special Reserve Mixed Media
SR-S  Special Reserve Script
SR-L  Special Reserve Locked
FH, FP  [Reserve] Foreign-Language H, P
FPR  [Reserve] Foreign-Language Perry Rhodan
C/RAN-H, C/RAN-P  Circulating Random H, P (non-fiction)
A  [Reserve] Art Book
B  [Reserve] Booklet
BLB  [Reserve] Big (or Better) Little Book
CAL  [Reserve] Calendar
CX, SCX  [Reserve] Comics, Small Comics
DICT  [Reserve] Dictionary
WRD  [Reserve] Weird Categories

Unearthed Errors

Found an error in the catalog? We’d like to know so we can fix it. Please give author and title as currently listed and what’s wrong (something is misspelled, the book should be listed as being in series X, etc).